OpenStudy (anonymous):

]If a sample of pure iodine contains 4.69 x 1022 atoms of iodine, how many moles of iodine are in the sample? 0.08 mol 0.13 mol 7.78 mol 12.84 mol

OpenStudy (anonymous):

thank you :)

OpenStudy (abb0t):

Use avogadro's number to convert. Avogadro's number is: \(6.022 \times 10^{23}\frac{ mol }{ atom }\)

OpenStudy (anonymous):

answer please

OpenStudy (anonymous):

????????

OpenStudy (abb0t):

You have to try figuring it out yourself first. I gave you the units. What you're doing is just like algebra, cancelling out units HINT: divide

OpenStudy (abb0t):

\(4.69 \times 10^{22} \frac{ \cancel{atoms} }{ 1 } \times \frac{ mol }{ 6.022 \times 10^{23}\cancel{atoms} }=\) ??